OpenStudy (anonymous):

If it takes two people six hours to move a rock, How long would it take three people to move the exact same rock?

OpenStudy (anonymous):

is there more information... it seems like it would depend on the strength of the person

OpenStudy (anonymous):

Are we assuming that these people all have the same strength?

OpenStudy (anonymous):

there is no more information given.... assume they all have the same strength

OpenStudy (anonymous):

It's a ratio then. 3 people to 2 people is x hours to 6 hours: 3/2 = x/6. Do the algebra, you get 3x = 12, which means x (the number of hours it will take with 3 people) = 4.

OpenStudy (anonymous):

four hours... i think (2ppl)(6hrs) = (3ppl)(xhrs) x hrs=4